OpenStudy (anonymous):

I dont know how to factor out GCF of a polynomial and the problem is 6y^2 + 12xy^2

OpenStudy (anonymous):

6y^2 = 2*3*y^2 12xy^2 = 2*2*3*y^2 so common factors are 2*3*y^2 i.e. 6y^2

OpenStudy (anonymous):

ans is 6y^2(1+2x)

OpenStudy (anonymous):

????????????